OpenStudy (anonymous):

12x^4 + 18x^3 + 3, does anyone know how to do this?

OpenStudy (anonymous):

you want to factorise it????

OpenStudy (anonymous):

its factoring out the GCF in polynomials

OpenStudy (anonymous):

GCF is 3

OpenStudy (anonymous):

because 12x^4 = 2*2*3*x^4 18x^3 = 2*3*3*x^3 3 = 3 So only 3 is common to all three terms

OpenStudy (anonymous):

Pls click the GOOD ANSWER button to show appreciation and to mark the question as solved and closed :))

OpenStudy (anonymous):

but i have multiple choice and the answers are a. 3x(4x^3 + 6x^2 + 1) b. 3(4x^4 + 6x^2) c. x(12x^3 + 18x^3 + 3) d. 3(4x^4 + 6x^3 + 1)

OpenStudy (anonymous):

answer is d) 3(4x^4 + 6x^3 + 1)

OpenStudy (anonymous):

thank you so much! this like litterally saved my life

OpenStudy (anonymous):

see you follow the method i hv shown after you separate out the GCF, you then see what is left for each one of the terms - this should come in the bracket

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

i hv answered the other question too..... pls check

OpenStudy (anonymous):

idk how i just signed up for this thing today

OpenStudy (anonymous):

u r welcome to the gang !!

OpenStudy (anonymous):

Thanks!! but i have problems left to do on this test and they are worth so many pionts and if i fail getting these wrong then im screwed, i fail the whole test

OpenStudy (anonymous):

u solve n post for us to confirm if your answer is ok or not That way you will learn also and if u make a mistake we can correct it :))

OpenStudy (anonymous):

but idk how to do the problem....i'll try and figure it out but i dont have much time

OpenStudy (anonymous):

give me a problem and ill explain better if you have a SKYPE account, we could have a voice chat and i can quickly teach you the method to follow

OpenStudy (anonymous):

i dont have skype

OpenStudy (anonymous):

ok then give a question n ill explain here

OpenStudy (anonymous):

Find the roots by applying the zero-product theorem 33. 3(x + 2)(x – 5) = 0 34. x(3x – 6) = 0 this is what im doing

OpenStudy (anonymous):

see once you put a value for x then (x+2) and (x-5) will reduce to a single number say A and B so 3(x+2)(x-5) =0 become 3*A*B = 0 This product can only become zero if either A is zero or B is zero. Understood till now ???

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

see the basic thing is that if you are multiplying even ten numbers and one of them is zero the entire product is zero so if A = 0 , this means that x+2=0 => x = -2 if B = 0, this means that x-5 = 0 => x = 5 so the two roots are x = -2 and x = 5

OpenStudy (anonymous):

Got it ?????

OpenStudy (anonymous):

If yes, you show how the second one is to be done and I'll say correct or not...

OpenStudy (anonymous):

no i dont get it, im sorry, im probably wasting your time

OpenStudy (anonymous):

i wish u had a SKYPE account... Actually it takes hardly 10 minutes to open an account and install skype on your computer. If u cud do it, i cud explain it to you better.....

OpenStudy (anonymous):

my parents said im not allowed to have skype even tho im 15, i wish i did too tho it would be alot better

OpenStudy (anonymous):

well............. it cud always b used under supervision........